# DrVoice: Parallel Speech-Text Voice Conversation Model via Dual-Resolution Speech Representations

### v3

@@ Appendix D Limitations and Future Work
Our future work will address limitations of this work and advance three key areas: (1) Enhancing Speech Generation Quality: The model currently exhibits weaker alignment between text and speech in speech-to-speech generation when compared to Qwen2.5-Omni. This performance gap can be attributed to a key architectural difference in the final speech generation stage. The Qwen2.5-Omni method directly feeds text into its talker module to ensure high textual fidelity. In contrast, DrVoice only inputs hidden states into its speech refined head (SRH). Therefore, a promising direction for future work is to incorporate text as an additional input to the SRH. It is hypothesized that this modification will provide more explicit textual guidance, further enhancing speech-text alignment and leading to a significant reduction in ASR-WER. (2) Enabling Full-Duplex Interaction: A crucial direction is to enable full-duplex interaction for more natural conversations. Inspired by recent advancements like Parrot (Wang et al., 2025), future work will investigate the use of a time-division multiplexing (TDM) input stream. This would permit DrVoice to accept user speech inputs during its own speech generation phase, thereby efficiently utilizing idle time slots and allowing for responsive, interruptible dialogue. (3) Expanding to General Audio and Multimodality: Finally, a promising avenue is to extend the model’s capabilities beyond speech-centric tasks to broader audio comprehension and synthesis, including the recognition and generation of music and environmental sounds. The subsequent integration of the visual modality will be a critical step toward developing a more comprehensive, multimodal conversational AI.

### v4

@@ Appendix C More Ablation and Analyses
SRH-Pretraining Policy. Two pre-training strategies were employed for the Speech Refined Head (SRH). The first strategy involves independently training a Text-to-Speech (TTS) task on the Qwen2.5-0.5B model in a T2M format with 25Hz token rate; the resulting trained weights are then loaded onto the SRH. In contrast, the second strategy freezes the Speech Encoder, Adaptor, Shared LLM Layer, and the Text Head, while exclusively training the SRH with a streaming TTS objective. Experiments demonstrated that the second strategy yields a substantial improvement in Speech-to-Speech (S2S) performance on the UltraEval-Audio benchmark, with the score increasing from 47.66 to 56.66. This significant gain is attributed to the closer alignment between the pre-training task in the second strategy and the format of the downstream task.
Reinforcement learning. To further enhance the model’s capabilities in real-world speech comprehension, instruction following, and reasoning, reinforcement learning (RL) methods were employed. Specifically, to address the scarcity of authentic speech data, a preference dataset was constructed using open-source ASR datasets. A text LLM was used to identify samples suitable for user queries, for which canonical responses were then generated by another text LLM. Subsequently, the DrVoice-SFT model performed inference on the original speech queries to establish the preference data. For the instruction-following capability, the open-source preference dataset  [^4] https://huggingface.co/datasets/allenai/tulu-3-pref-personas-instruction-following was utilized. Training for these first two objectives was performed using the DPO (DBLP:conf/nips/RafailovSMMEF23) algorithm. Finally, to improve reasoning ability, the model underwent GSPO (DBLP:journals/corr/abs-2507-18071) training on mathematical problems  [^5] https://github.com/openai/prm800k.
@@ Appendix D Limitations and Future Work
Our future work will address limitations of this work and advance two key areas: (1) Enabling Full-Duplex Interaction: A crucial direction is to enable full-duplex interaction for more natural conversations. Inspired by recent advancements like Parrot (wang2025parrot), future work will investigate the use of a time-division multiplexing (TDM) input stream. This would permit DrVoice to accept user speech inputs during its own speech generation phase, thereby efficiently utilizing idle time slots and allowing for responsive, interruptible dialogue. (2) Expanding to General Audio and Multimodality: Finally, a promising avenue is to extend the model’s capabilities beyond speech-centric tasks to broader audio comprehension and synthesis, including the recognition and generation of music and environmental sounds. The subsequent integration of the visual modality will be a critical step toward developing a more comprehensive, multimodal conversational AI.